# Bias in Large Language Models: Origin, Evaluation, and Mitigation

### v1

@@ 4.4 Human-Involved Bias Evaluation Methods
Human-involved evaluation incorporates human judgment to identify complex and context-dependent biases in LLM outputs.
@@ 5 Bias Mitigation
Biases are often present in unstructured data, and LLMs trained on such data can not only learn these biases but sometimes amplify them as well (Bolukbasi et al., 2016; Sheng et al., 2019). Various methods of bias mitigation exist under different contexts, such as social bias (Garimella et al., 2021, 2022), political bias (Liu et al., 2022), and stereotype bias (Nie et al., 2024b). Instead of reviewing these methods for bias mitigation by bias contexts, we categorize them by the timing of debiasing. More specifically, methods for mitigating biases can be categorized into three distinct approaches: (1) preprocessing the input data prior to training LLMs; (2) changing the model’s architecture and/or picking the best model based on the predefined bias evaluation metrics; (3) improving unbiasedness by calibrating the outputs at the decision-making stage. We call these three approaches as pre-model debiasing, intra-model debiasing, and post-model debiasing, respectively. Table 4 provides an overview of the advantages and disadvantages on each approach.
@@
Pre-model Resampling; Data augmentation; Expert intervention Time-saving without model training; Time-consuming to annotate bias cases, Ineffectiveness with a self-biased model, Privacy issue of resampling
Intra-model Equalization and declustering; Movement pruning; Transfer learning; Dropout regularization; Causal inference Flexibility in mitigating various types of biases, Strong performance in empirical studies; Time-consuming for modifying and training models
Post-model Reinforced calibration; Self-Debias; Projection-based methods; Causal prompting Time-saving without model training; Demand of a large amount of data
@@ 5.1 Pre-model debiasing
Pre-model debiasing approach involves preprocessing the input data before training LLMs so that the entire training data are balanced and representative of the entire population. Typical ways of balancing training data include oversampling, undersampling, and synthetic data generation (Ferrara, 2023). We review the commonly used pre-model debiasing methods below.
One of the widely used methods is counterfactual data augmentation (CDA), which is often used to mitigate bias in the literature (Zmigrod et al., 2019; Webster et al., 2020; Barikeri et al., 2021; Mondal and Lipizzi, 2024). Counterfactual Data Substitution (CDS), a variant of CDA, is also proposed to randomly substitute potentially biased text to avoid duplication (Maudslay et al., 2019). Generally speaking, CDA rebalances a corpus by sawpping bias attribute words in a dataset. For example, we may swap the sentence
"The professor came to the classroom but he forgot to bring his laptop"
to
@@
Meade et al. (2022) further empirically evaluates CDA for mitigating religious bias by swapping religious terms.
@@
Serouis and Sèdes (2024) proposes an iterative method that human can oversee the preprocessing step. That is, before training begins, a dataset report is sent to a LLM. An expert then decides the next step: either issuing a command to the LLM to create new training data or starting the training process. The decision is made by measuring the quality of generated data. If a command is used to generate alternative training data, iterative feedback between the expert and the LLM continues until predefined criteria are satisfied, leading to a decision to proceed with model training. After training and experimentation, detailed results concerning protected groups are provided to the LLM for summarization (particularly for multiple groups and subgroups) or directly to the expert. Based on these comprehensive reports, the expert may choose to use the LLM to produce a revised dataset version to mitigate potential biases.
5.2 Intra-mdoel debiasing
@@
Intra-mdoel debiasing works by selecting the model that prioritizes unbiasedness and fairness during the model fitting. For example, Garimella et al. (2021) proposes a method to mitigate explicit and implicit biases in BERT using existing and newly proposed loss functions. The proposed work focuses on gender bias and includes a list of target-defined word pairs and the text dataset as input. It takes the pretrained BERT and further trains it on the given input, while mitigating the existing social biases by equalizing and declustering. The equalizing stage aims to equalize the associations of neutral words by defining a equalizing loss (Qian et al., 2019). Observing that words are stereotypically associated to a group (e.g., delicate, pink, beautiful), authors develop a declustering stage to decluster the implicit clusters by defining a declustering loss. Similar technique is also applied to mitigate social stereotypes when exposed to different demographic groups (Garimella et al., 2022). Moreover, Joniak and Aizawa (2022) presents an innovative framework for examining bias in pre-trained transformer-based language models through movement pruning. It identifies a subset of the model that contains less bias than the original by pruning it while fine-tuning on a debiasing objective. It explores pruning attention heads, which are crucial components of transformers, by pruning square blocks and introducing a novel method for pruning entire heads. An interesting bias-performance trade-off about gender bias from this study is: the level of bias increases as model’s performance improves.
@@
Another line of works focuses on the technique of transfer learning (Zhuang et al., 2020; Azunre, 2021; Ge et al., 2024; Liu, 2024). Transfer learning leverages knowledge learned from a pre-trained model to improve the performance of a new model on a relevant task but with fewer data points. More specifically, it fine-tunes a pre-trained model on the new task by updating the weights of neurons or modifying the network architecture. For example, FairDistillation (Delobelle and Berendt, 2022), a cross-lingual method based on knowledge distillation, is used to build smaller language models while managing specific biases. It leverages existing language models as a teacher, providing a richer training signal without the need for retraining from scratch. To avoid transferring learned correlations to new language models, it further substitutes CDA’s augmentation strategy with probabilistic rules between tokens. Furthermore, we notice that a few works utilize multilinguality as bias mitigation (Ahn and Oh, 2021; Levy et al., 2023; Nie et al., 2024a). These works advocate multilingual training to mitigate biases. They demonstrate the effectiveness of multilinguality in reducing bias and boosting prediction accuracy (Nie et al., 2024a).
@@
Another useful method is dropout regularization (Srivastava et al., 2014; Webster et al., 2020). Briefly, they explore the impact of increasing dropout parameters for BERT and ALBERT’s attention weights and hidden activations, followed by an additional phase of pre-training. Experiments reveal that dropout regularization reduces gender bias in these models. They propose that dropout disrupts the attention mechanisms in BERT and ALBERT, thereby preventing the models from learning undesirable word associations.
@@
Grounded in the theory of causal inference, causality-guided methods are investigated due to the theoretical guarantees and good generalization (Da et al., 2024; Cai et al., 2024). For instance, researchers propose a bias reduction method for sentiment classification (Da et al., 2024). This framework treats the neural network as a causal Directed Acyclic Graph (DAG). Borrowing the concept of causal mediation analysis for neural network (Vig et al., 2020), authors study the direct and indirect effects of individual neurons to identify parts of a model where gender bias is causally implicated. Finally, authors apply counterfactual training to only fine-tune the BERT target layers. Moreover, Cai et al. (2024) uses causal mediation analysis to trace the causal effects of different components’ activation within a LLM. In this way, we are able to investigate the storage locations and generation mechanisms of gender bias in LLMs. Least Square Debias Method (LSDM) is proposed to modify parameters to mitigate gender bias in models.
@@ 5.3 Post-model debiasing
Post-model debiasing mitigates bias by adjusting the output of LLMs. The benefits of post-model debiasing are no changes in the training data and no changes in the architecture of LLMs. For example, reinforced calibration can be used to mitigates political bias in GPT-2 (Liu et al., 2021, 2022). It keeps the architecture of GPT-2 unchanged but calibrates bias during the generation. More specifically, authors present a reinforcement learning based framework for political biases in two modes: word embedding debias and classifier guided debias. For the word embedding debias, authors pick the unbiased token at each generation step by forcing neutral words to maintain equal distances from groups of sensitive words (e.g., male and female) in the embedding space (Bolukbasi et al., 2016; Park et al., 2018; Zhao et al., 2018c). The classifier guided debiasing works when the bias is not purely word level (Bordia and Bowman, 2019).
@@
Self-Debias is another work that leverages model’s internal knowledge to generate unbiased prompts (Schick et al., 2021). The necessary condition for this approach is that if a model is able to discern the presence of undesired biases, it should be able to avoid them at test time. Self-Ddebias first detects when the models output undesirable attributes based on internal knowledge. It then reduces the probability of producing biased test (such as sexist, racist, and violent) by a decoding algorithm. This has been empirically evaluated as a strong debiasing technique (Meade et al., 2022).
@@
Projection-based methods are also shown to be useful for mitigating biases. Notable methods such as SENT-DEBIAS (Liang et al., 2020a) and INLP (Ravfogel et al., 2020) work by projecting model’s representations onto a subspace. More specifically, SENT-DEBIAS involves four steps: (1) identifying words that display biased attributes (e.g., he and she); (2) embedding these words within biased attribute sentences to generate their sentence representations by CDA; (3) estimating the bias subspace within these sentence representations by principle component analysis (PCA); (4) debiasing general sentences by removing their projection onto this bias subspace. INLP roughly debiases a model’s representations by training a linear classifier to predict the protected attribute we want to remove (e.g., gender) from the representations. The representations are then debiased by projecting them into the nullspace of the trained classifier’s weight matrix, effectively eliminating all the information the classifier used to predict the protected attribute from the representation. This process can be applied iteratively to further debias the representation.
@@
In alignment with causal methods in intra-model debiasing, researchers also develop methods based on causal prompting (Li et al., 2024a; Zhang et al., 2024). Prompting methods gain popularity since the general public has no access to the model’s internal structure due to business interest. Specifically, Li et al. (2024a) proposes causal prompting based on front-door adjustment (Pearl et al., 2016). The proposed method modifies prompts without access to the parameters and logits of LLMs. First, it queries LLMs to generate chain-of-thoughts (CoTs) $m$ times with the input prompt (demonstration examples and a question of the test example). An encoder-based clustering algorithm is applied to these CoTs and top $K$ representative CoTs are selected. Next, it retrieves the optimal demonstration examples for each representative. Finally, LLMs are queried $T$ times to obtain $T$ answers for each representative CoT, and the final answer is obtained by a weighted voting. Zhang et al. (2024) constructs detailed causal models for both the underlying data-generating process of the training corpus and the LLM’s reasoning process. They recognize that selection mechanisms are crucial in determining how the LLM’s output can be influenced by different prompts. Additionally, they propose a causality-guided debiasing framework. By gaining causal insights into the data-generating processes, the necessary conditions that prompt-design strategies must meet to achieve principled and effective debiasing are identified.
@@ 6 Ethical Concerns and Legal Challenges
Due to rapid advancements and successful integration of LLMs into systems that are widespread throughout our society, more and more are raising concerns that LLMs can learn and amplify societal biases (gender, race, religion, age, etc.) embedded in the enormous Internet data they are trained on (Gallegos et al., 2024). As a result, LLMs can potentially generate biased outputs among different social groups, which falls under the ethical concerns of AI. Furthermore, disparate treatments led by biased outputs in decision-making processes can result in legal challenges under anti-discrimination laws.
We categorize biased outputs and disparate treatments according to a pre-established taxonomy: representational harms and allocational harms (Blodgett et al., 2020). Representational harms refers to when an automated system exhibits negative altitudes or misrepresents, even ignores existence of certain social groups. Allocational harms refers to when an automated system allocate resources in favor of some social groups over the others.
After reviewing many types of representational harms described in the literature, we constrained our discussion to four categories that can cover most representational harms / ethical concerns:
@@
    • Result 1: Reinforced social bias against certain social groups,
@@
    • Result 1: Reinforced normativity of the dominant social group and implicit exclusion or devaluation or other groups.
@@
    • Result 1: Allocational harms
Stereotyping. LLMs can reinforce over-simplified and often negative labels on certain social groups. Abid et al. (2021) demonstrate that LLMs captures Muslim-violence bias in various uses cases and the Muslin-violence bias is more severe than commonly-seen biases of other religious groups. Bender et al. (2021) describe how LLMs like referring to "women doctors" since even in today’s world, doctor is a male-dominant occupation. Stereotypes perpetuated and even exacerbated by LLMs can further translate into offensive languages that target a social group. Gehman et al. (2020) evaluate five language models (including GPT-1 to 3) using a "REALTOXICITYPROMPTS" machine and show that all five model degenerates into toxicity over 0.5 (assigned by a widely-accepted toxicity classifier) within less than 100 generations from non-toxic prompts. Toxic generations from language models can range from the like of "Islamic slaughtering" to "lying white media".
Misrepresentation. The disproportionate representation of marginalized or minority social groups in LLM training data can lead to implicit exclusion of marginalized groups and mischaracterization of certain groups. Bender et al. (2021) raise an example of ‘both genders’ that exclude non-binary identities from the context. Navigli et al. (2023) provide several examples of misrepresentation across gender, age, nationality, ethnicity, etc. For example, "they are not very good at English" as a response to "they are Chinese" or "she is ignorant" as a response to "she is White".
Disparate system performance. Koenecke et al. (2020) reveal that all of the five state-of-the-art LLM-based automated speech recognition (ASR) systems show higher word error rates for black speakers than white speakers. Saunders and Byrne (2020) demonstrate that the over-prevalence of male-centric forms in Machine Translation training data result in better translations for male-centric sentences and for sentences that contains male-dominant roles such as doctors and developers. Next, we discuss disparate system performance in LLM-automated hiring and healthcare tools, which often lead to allocational harms that raise anti-discrimination legal challenges.
@@
    • LLM-based Hiring Tools: Resume and Cover Letter Screening
    • LLM-based Health Care Tools: AI Diagnostics
Allocational harms. Social biases preserved by LLMs would likely lead to direct or indirect discrimination actions from LLMs. Ferrara (2023) argues that AI-automated tools may reserve and exacerbate societal biases embedded in the training data. LLM-aided resume analyzing tools may disqualify people based on non-traditional backgrounds instead of job-related qualifications. In 2018, Amazon reportedly took an internal AI recruiting tool off the shelf after being accused of bias against women since the resumes in the training data are predominantly from white males. Another example is that LLM-aided healthcare tools may generate inaccurate predictions and treatment recommendations for under-represented groups, potentially influencing healthcare decision-making and resulting in unfair medical resource allocation. Zack et al. (2024) discovered that GPT-4 is less likely to recommend more expensive diagnostician procedures to Black people and more likely to predict Hispanic women are more likely to hide alcohol abusing history than Asian women.
7 Conclusion

### v2

@@ 4.4 Human-Involved Bias Evaluation Methods
Human-involved evaluation incorporates human judgment to identify complex and context-dependent biases in LLM outputs. Human-involved evaluation is defined by the source of judgment (i.e., human assessors), whereas domain-specific evaluation is defined by application-specific norms, risks, and fairness criteria (e.g., healthcare, law, education); accordingly, a domain-specific evaluation may be human-involved, automated, or both.
@@
To reduce subjectivity in human-involved bias evaluation, the evaluation protocol (Figure 4) should be specified explicitly rather than left implicit. At a minimum, such a protocol should define: (1) the target population of evaluators and the rationale for their selection; (2) the diversity requirements of the evaluator pool, including demographic, cultural, linguistic, and professional diversity when the task involves socially situated judgments; (3) whether domain experts, affected-group members, crowd workers, or mixed panels are used; (4) standardized annotation guidelines, examples, and rating rubrics; and (5) the procedure for pilot calibration before formal evaluation. In practice, evaluators should assess outputs independently and, when possible, without access to model identity or study hypotheses, in order to reduce anchoring and confirmation effects. The manuscript should also report the number of evaluators, their relevant background characteristics, the assignment protocol, and inter-annotator agreement statistics when applicable Artstein (2017). For disagreement handling, a pre-specified rule should be provided, such as majority vote for straightforward labels, adjudication by a trained senior reviewer, or discussion-based consensus for context-sensitive cases; unresolved disagreements may also be retained and analyzed separately as indicators of ambiguity in the bias construct itself. Such protocol design helps ensure that human evaluation is systematic, transparent, and reproducible, rather than relying solely on subjective impressions.
Figure 4: Human evaluation protocol design.
@@ 5 Bias Mitigation
Biases are often present in unstructured data, and LLMs trained on such data can not only learn these biases but sometimes amplify them as well (Bolukbasi et al., 2016; Sheng et al., 2019). Various methods of bias mitigation exist under different contexts, such as social bias (Garimella et al., 2021, 2022), political bias (Liu et al., 2022), and stereotype bias (Nie et al., 2024). Instead of reviewing these methods for bias mitigation by bias contexts, we categorize them by the timing of debiasing. More specifically, methods for mitigating biases can be categorized into three distinct approaches: (1) preprocessing the input data prior to training LLMs; (2) changing the model’s architecture and/or picking the best model based on the predefined bias evaluation metrics; and (3) improving unbiasedness by calibrating the outputs at the decision-making stage. We call these three approaches pre-model debiasing, intra-model debiasing, and post-model debiasing, respectively. Table 4 provides an overview of the advantages and disadvantages of each approach.
@@
Pre-model Resampling; data augmentation; expert intervention Time saving without model training; time consuming to annotate bias cases, ineffectiveness with a self-biased model, privacy issue of resampling
Intra-model Equalization and declustering; movement pruning; transfer learning; dropout regularization; causal inference Flexibility in mitigating various types of biases, strong performance in empirical studies; time consuming for modifying and training models
Post-model Reinforced calibration; Self-Debias; projection-based methods; causal prompting Time saving without model training; demand of a large amount of data
Figure 5 provides a visual summary of the three-stage debiasing pipeline and the representative techniques within each stage.
Pre-Model Debiasing Intra-Model Debiasing Post-Model Debiasing TrainingInference CDA/CDS Resampling Expert curation Equalization loss Movement pruning Transfer learning Dropout/Causal Self-Debias Projection (INLP) RL calibration Causal prompting
Figure 5: Overview of the three-stage bias mitigation pipeline. Pre-model methods intervene on training data, intra-model methods modify the training process or model architecture, and post-model methods adjust outputs at inference time. Representative techniques are listed under each stage.
@@ 5.1 Pre-Model Debiasing
Pre-model debiasing approaches involve preprocessing the input data before training LLMs, so that the entire training data are balanced and representative of the entire population. Typical ways of balancing training data include oversampling, undersampling, and synthetic data generation (Ferrara, 2023). We review the commonly used pre-model debiasing methods below.
One of the widely used methods is counterfactual data augmentation (CDA), which is often used to mitigate bias in the literature (Zmigrod et al., 2019; Webster et al., 2020; Barikeri et al., 2021; Mondal and Lipizzi, 2024). Counterfactual data substitution (CDS), a variant of CDA, is also proposed to randomly substitute potentially biased text to avoid duplication (Maudslay et al., 2019). Generally speaking, CDA rebalances a corpus by swapping bias attribute words in a dataset. For example, we may swap the sentence
“The professor came to the classroom but he forgot to bring his laptop”
with
@@
Meade et al. Meade et al. (2022) further empirically evaluate CDA for mitigating religious bias by swapping religious terms.
@@
Serouis et al. Serouis and Sèdes (2024) propose an iterative method in which a human can oversee the preprocessing step. That is, before training begins, a dataset report is sent to an LLM. An expert then decides the next step: either issuing a command to the LLM to create new training data or starting the training process. The decision is made by measuring the quality of the generated data. If a command is used to generate alternative training data, iterative feedback between the expert and the LLM continues until predefined criteria are satisfied, leading to a decision to proceed with model training. After training and experimentation, detailed results concerning protected groups are provided to the LLM for summarization (particularly for multiple groups and subgroups) or directly to the expert. Based on these comprehensive reports, the expert may choose to use the LLM to produce a revised dataset version to mitigate potential biases.
5.2 Intra-Model Debiasing
@@
Intra-model debiasing works by selecting the model that prioritizes unbiasedness and fairness during the model fitting. For example, Ref. Garimella et al. (2021) proposes a method to mitigate explicit and implicit biases in BERT using existing and newly proposed loss functions. The proposed work focuses on gender bias and uses a predefined list of target word pairs and a text dataset as inputs. It takes the pretrained BERT and further trains it on the given input, while mitigating the existing social biases by equalizing and declustering. The equalizing stage aims to equalize the associations of neutral words by defining an equalizing loss (Qian et al., 2019). Observing that words are stereotypically associated with a group (e.g., delicate, pink, beautiful), the authors develop a declustering stage to decluster the implicit clusters by defining a declustering loss. A similar technique is also applied to mitigate social stereotypes when exposed to different demographic groups (Garimella et al., 2022). Moreover, Ref. Joniak and Aizawa (2022) presents an innovative framework for examining bias in pretrained transformer-based language models through movement pruning. It identifies a subset of the model that contains less bias than the original by pruning it while fine-tuning on a debiasing objective. It explores pruning attention heads, which are crucial components of transformers, by pruning square blocks and introducing a novel method for pruning entire heads. An interesting bias–performance trade-off for gender bias from this study is that the level of bias increases as the model’s performance improves.
@@
Another line of studies focuses on the technique of transfer learning (Zhuang et al., 2020; Azunre, 2021; Ge et al., 2023; Liu, 2024). Transfer learning leverages knowledge learned from a pretrained model to improve the performance of a new model on a relevant task but with fewer data points. More specifically, it fine-tunes a pretrained model on the new task by updating the weights of neurons or modifying the network architecture. For example, FairDistillation (Delobelle and Berendt, 2022), a cross-lingual method based on knowledge distillation, is used to build smaller language models while managing specific biases. It leverages existing language models as a teacher, providing a richer training signal without the need for retraining from scratch. To avoid transferring learned correlations to new language models, it further substitutes CDA’s augmentation strategy with probabilistic rules between tokens. Furthermore, we notice that a few works utilize multilinguality as bias mitigation (Ahn and Oh, 2021; Levy et al., 2023; Nie et al., 2024). These works advocate multilingual training to mitigate biases. They demonstrate the effectiveness of multilinguality in reducing bias and boosting prediction accuracy (Nie et al., 2024).
@@
Another useful method is dropout regularization, which is studied in (Srivastava et al., 2014; Webster et al., 2020). Briefly, they explore the impact of increasing dropout parameters for BERT’s and ALBERT’s attention weights and hidden activations, followed by an additional phase of pretraining. Their experiments reveal that dropout regularization reduces gender bias in these models. They propose that dropout disrupts the attention mechanisms in BERT and ALBERT, thereby preventing the models from learning undesirable word associations.
@@
Grounded in the theory of causal inference, causality-guided methods have been investigated due to the theoretical guarantees and good generalization (Da et al., 2024; Cai et al., 2024). For instance, researchers have proposed a bias reduction method for sentiment classification (Da et al., 2024). This framework treats the neural network as a causal directed acyclic graph (DAG). Borrowing the concept of causal mediation analysis for neural networks (Vig et al., 2020), the authors study the direct and indirect effects of individual neurons to identify parts of a model where gender bias is causally implicated. Finally, the authors apply counterfactual training to only fine-tune the BERT target layers. Moreover, Ref. Cai et al. (2024) uses causal mediation analysis to trace the causal effects of different components’ activation within an LLM. In this way, they were able to investigate the storage locations and generation mechanisms of gender bias in LLMs. The least square debias method (LSDM) was proposed to modify parameters to mitigate gender bias in models.
@@ 5.3 Post-Model Debiasing
Post-model debiasing mitigates bias by adjusting the output of LLMs. The benefits of post-model debiasing are that there are no changes in the training data and no changes in the architecture of the LLM. For example, reinforced calibration can be used to mitigate political bias in GPT-2 (Liu et al., 2021, 2022). It keeps the architecture of GPT-2 unchanged but calibrates bias during the generation. More specifically, researchers have presented a reinforcement learning-based framework for political biases in two modes: word embedding debias and classifier-guided debias. For the word embedding debias, the researchers pick the unbiased token at each generation step by forcing neutral words to maintain equal distances from groups of sensitive words (e.g., male and female) in the embedding space (Bolukbasi et al., 2016; Park et al., 2018; Zhao et al., 2018). The classifier-guided debiasing works when the bias is not purely word-level (Bordia and Bowman, 2019).
@@
Self-Debias is another work that leverages the model’s internal knowledge to generate unbiased prompts (Schick et al., 2021). The necessary condition for this approach is that if a model is able to discern the presence of undesired biases, it should be able to avoid them at test time. Self-Debias first detects when the models output undesirable attributes based on internal knowledge. It then reduces the probability of producing a biased test (such as sexist, racist, and violent) using a decoding algorithm. This has been empirically evaluated as a strong debiasing technique (Meade et al., 2022).
@@
Projection-based methods have also been shown to be useful for mitigating biases. Notable methods such as SENT-DEBIAS (Liang et al., 2020) and INLP (Ravfogel et al., 2020) work by projecting a model’s representations onto a subspace. More specifically, SENT-DEBIAS involves four steps: (1) identifying words that display biased attributes (e.g., he and she); (2) embedding these words within biased attribute sentences to generate their sentence representations by CDA; (3) estimating the bias subspace within these sentence representations by principle component analysis (PCA); and (4) debiasing general sentences by removing their projection onto this bias subspace. INLP roughly debiases a model’s representations by training a linear classifier to predict the protected attribute we want to remove (e.g., gender) from the representations. The representations are then debiased by projecting them into the nullspace of the trained classifier’s weight matrix, effectively eliminating all the information the classifier used to predict the protected attribute from the representation. This process can be applied iteratively to further debias the representation.
@@
In alignment with causal methods in intra-model debiasing, researchers have also developed methods based on causal prompting (Li et al., 2024; Zhang et al., 2024). Prompting methods have gained in popularity since the general public has no access to the model’s internal structure due to business interests. Specifically, Ref. Li et al. (2024) proposes causal prompting based on front-door adjustment (Pearl et al., 2016). The proposed method modifies prompts without access to the parameters and logits of LLMs. First, it queries LLMs to generate chains of thought (CoTs) $m$ times with the input prompt (demonstration examples and a question of the test example). An encoder-based clustering algorithm is applied to these CoTs and the top $K$ representative CoTs are selected. Next, it retrieves the optimal demonstration examples for each representative. Finally, the LLMs are queried $T$ times to obtain $T$ answers for each representative CoT, and the final answer is obtained by a weighted voting. Ref. Zhang et al. (2024) constructs detailed causal models for both the underlying data-generating process of the training corpus and the LLM’s reasoning process. They recognize that selection mechanisms are crucial in determining how the LLM’s output can be influenced by different prompts. Additionally, they propose a causality-guided debiasing framework. By gaining causal insights into the data-generating processes, the necessary conditions that prompt-design strategies must meet to achieve principled and effective debiasing are identified.
5.4 Critical Comparison of Mitigation Strategies
Having reviewed the three categories of debiasing approaches, we now offer a critical comparison of their effectiveness, computational requirements, and practical trade-offs. Table LABEL:tab:mitigation_comparison provides an estimated comparison of the computational costs across representative methods.
Table 5: Estimated computational cost comparison of representative bias mitigation strategies. Costs are approximate and depend on model size and implementation.
Category Method Key Mechanism Relative Cost GPU Requirement
Pre-model CDA/CDS Augmenting training data Low Minimal
Pre-model Expert curation Human review of data Medium None
Intra-model Equalization loss Modified training objective High Full training
Intra-model Movement pruning Pruning biased subnetworks High Full training
Intra-model FairDistillation Knowledge distillation Medium–high Distillation run
Intra-model Dropout regularization Additional pretraining Medium Pretraining pass
Post-model Self-Debias Modified decoding Low Inference only
Post-model SENT-DEBIAS/INLP Subspace projection Low–medium Inference + PCA
Post-model Reinforced calibration RL-based generation Medium RL training
Post-model Causal prompting Prompt modification Low–medium Multiple queries
Several key trade-offs emerge from this comparison. Pre-model methods are conceptually straightforward and computationally inexpensive but cannot address biases that arise from model architecture or training dynamics. They are also limited in their ability to handle biases that are not explicitly encoded in individual words or phrases but emerge from complex distributional patterns. Intra-model methods offer the most direct intervention into the bias-producing mechanisms but require full model retraining or fine-tuning, which is computationally expensive and may not be feasible for the largest contemporary LLMs. Furthermore, intra-model methods are model-specific and may not generalize across architectures. Post-model methods are attractive for their flexibility and low computational overhead, as they do not require modifying the training data or model parameters. However, they address only the symptoms of bias rather than its root causes, and their effectiveness depends on the ability to detect bias at inference time. Self-Debias, for instance, relies on the assumption that the model can recognize its own biases, which may not hold for all bias types, especially subtle or intersectional biases.
In practice, a combination of methods across all three stages is likely to yield the most robust debiasing outcomes. Pre-model data curation can reduce the most egregious data-level biases, intra-model techniques can address biases encoded in model representations, and post-model methods can serve as a final safety net during deployment. Future research should focus on developing integrated debiasing pipelines that coordinate interventions across these stages.
@@ 6 Ethical Concerns and Legal Challenges
Due to rapid advancements and successful integration of LLMs into systems that are widespread throughout our society, more and more concerns are being raised that LLMs can learn and amplify societal biases (gender, race, religion, age, etc.) embedded in the enormous Internet data they are trained on (Gallegos et al., 2024). As a result, LLMs can potentially generate biased outputs among different social groups, which falls under the ethical concerns of AI. Furthermore, disparate treatment caused by biased outputs in decision-making processes can result in legal challenges under anti-discrimination laws.
We categorize biased outputs and disparate treatment according to a pre-established taxonomy: representational harms and allocational harms (Blodgett et al., 2020). Representational harms refer to when an automated system exhibits negative attitudes or misrepresents, even ignores existence of certain social groups. Allocational harms refer to when an automated system allocate resources in favor of some social groups over others.
After reviewing many types of representational harms described in the literature, we constrained our discussion to four categories that can cover most representational harms or ethical concerns:
@@
    • Result 1: Reinforced social bias against certain social groups.
@@
    • Result 1: Reinforced normativity of the dominant social group and implicit exclusion or devaluation of other groups.
@@
    • Result 1: Allocational harms.
Stereotyping: LLMs can reinforce over-simplified and often negative labels attached to certain social groups. Abid et al. (2021) demonstrate that LLMs capture Muslim-violence bias in various use cases and the Muslim-violence bias is more severe than commonly seen biases for other religious groups. Bender et al. (2021) describe how LLMs tend to refer to “women doctors” since even in today’s world, doctor is a male-dominant occupation. Stereotypes perpetuated and even exacerbated by LLMs can further translate into offensive language that targets a social group. Gehman et al. (2020) evaluate five language models (including GPT-1 to 3) using a “REALTOXICITYPROMPTS” machine and show that all five models degenerate into toxicity over 0.5 (assigned by a widely accepted toxicity classifier) within less than 100 generations from non-toxic prompts. Toxic generations from language models can range from the likes of “Islamic slaughtering” to “lying white media”.
Misrepresentation: The disproportionate representation of marginalized or minority social groups in LLM training data can lead to implicit exclusion of marginalized groups and mischaracterization of certain groups. Bender et al. (2021) raise an example of ‘both genders’ that exclude non-binary identities from the context. Navigli et al. (2023) provide several examples of misrepresentation across gender, age, nationality, ethnicity, etc. For example, “they are not very good at English” as a response to “they are Chinese” or “she is ignorant” as a response to “she is White”.
Disparate system performance: Koenecke et al. (2020) reveal that all of the five state-of-the-art LLM-based automated speech recognition (ASR) systems show higher word error rates for black speakers than white speakers. Saunders and Byrne (2020) demonstrate that the over-prevalence of male-centric forms in machine translation training data result in better translations for male-centric sentences and for sentences that contain male-dominant roles such as doctors and developers. Next, we discuss disparate system performance in LLM-automated hiring and healthcare tools, which often leads to allocational harms that raise anti-discrimination legal challenges.
@@
    • LLM-based hiring tools: Resume and cover letter screening.
    • LLM-based healthcare tools: AI diagnostics.
Allocational harms: Social biases preserved by LLMs can translate into discriminatory allocation of opportunities and services—for example, by steering who gets shortlisted for jobs or which patients receive particular diagnostic workups—thereby creating disparate treatment or disparate impact in high-stakes settings.
In addition to ethical concerns, allocation harms intersect with legal and regulatory frameworks intended to prevent discrimination. In the U.S., federal employment discrimination protections apply to AI and other automated tools “just as they apply to other employment practices”, including recruitment and screening; employers remain responsible for ensuring that selection procedures do not unlawfully exclude protected groups, even when tools are provided by vendors EEOC (2023). At the subnational level, New York City’s Local Law 144 regulates “automated employment decision tools” by requiring recent bias audits, public summaries, and advance notice before use in hiring or promotion decisions New York City Council (2021). In healthcare, the 2024 Section 1557 final rule prohibits discrimination through the use of patient care decision support tools and requires covered entities to make reasonable efforts to identify high-risk uses and mitigate discrimination when such tools are used in clinical decision making U.S. Department of Health and Human Services (2024). In the EU, the AI Act adopts a risk-based approach: “high risk” includes recruitment tools (e.g., CV sorting) and AI-based medical software, with obligations such as risk management, data quality to reduce discriminatory outcomes, documentation, and human oversight European Commission (2023).
Despite these legal and regulatory frameworks, real-world examples demonstrate that allocation harms persist. Ferrara (2023) argues that AI-automated tools may preserve and exacerbate societal biases embedded in the training data. LLM-aided resume-analyzing tools may disqualify people based on non-traditional backgrounds instead of job-related qualifications. In 2018, Amazon reportedly took an internal AI recruiting tool off the shelf after being accused of bias against women since the resumes in the training data are predominantly from white males. Another example is that LLM-aided healthcare tools may generate inaccurate predictions and treatment recommendations for under-represented groups, potentially influencing healthcare decision making and resulting in unfair medical resource allocation. Zack et al. (2024) discovered that GPT-4 is less likely to recommend more expensive diagnostic procedures to Black people and more likely to predict Hispanic women are more likely to hide an alcohol-abuse history than Asian women.
7 Conclusions